1. Curriculum planners decided to determine whether a curriculum change should be
instituted to increase the reading comprehension of sixth-graders in Metro Manila.
Then mean reading comprehension score under the old curriculum was 92 on a
standard reading comprehension test. After the new curriculum was used for one
quarter, the mean of a sample of 400 students was 95. Assuming that the standard
deviation for the entire population of sixth-graders in Metro Manila is the same
under the new curriculum as under the old (s = 10), test at significance level 0.01
whether there would be an improvement in reading comprehension in general if
the entire population of sixth-graders in Metro Manila were taught using the new
curriculum.

Answer by stanbon(75887)


Curriculum planners decided to determine whether a curriculum change should be
instituted to increase the reading comprehension of sixth-graders in Metro Manila.
Then mean reading comprehension score under the old curriculum was 92 on a
standard reading comprehension test. After the new curriculum was used for one
quarter, the mean of a sample of 400 students was 95. Assuming that the standard
deviation for the entire population of sixth-graders in Metro Manila is the same
under the new curriculum as under the old (s = 10), test at significance level 0.01
whether there would be an improvement in reading comprehension in general if
the entire population of sixth-graders in Metro Manila were taught using the new
curriculum.
------------
Ho: u(new) <= 92
Ha: u(new) > 92 (claim)
-----
t(95) = (95-92)/[10/sqrt(400)] = 3/(10/20) = 6
p-value = P(t > 6 when df = 399) = tcdf(6,100,399) = 2.22x10^-9
----
Since the p-value is less than 1%, reject Ho.
The test results support the claim.
-----------
Cheers,
Stan H.